ADNAN MUHAMMED

Grade 12,

Javed was given 5 / 7 of a basket of oranges. What fraction of oranges was left in the basket?

Javed was given 5 / 7 of a basket of oranges. What fraction of oranges was left in the basket?

Grade:12

2 Answers

Priyanshu Singh
3 years ago
let the remaining fraction be x
5/7 + x = 1[ 1 is the whole]
x = 1 - 5/7
x = 2/7
therefore the remaining fraction of oranges = 2/7
Harshit Singh
3 years ago
Dear Student

Fraction of oranges given to Javed = 5 / 7
Fraction of oranges left in the basket = 1–5 / 7
=7 / 7–5 / 7
= (7–5) / 7
= 2/7

Thanks